Judy made 84% on her last math test. If she missed 4 problems, how many questions were on the test?

Answer by jim_thompson5910(35256)

Let x = number of questions on the test


She missed 4 questions, so she got x-4 questions correct.


84% of (number of questions on test) = number of correct answers


0.84x = x - 4


0.84x-x = -4


-0.16x = -4


x = -4/(-0.16)


x = 25


So there are 25 questions total on the test. She missed 4 of them, so she got 25-21 = 24 questions correct. Notice how 21/25 = 0.84 ---> 0.84*100 = 84%